OpenStudy (anonymous):

The surface area of the top surface of the water in a circular swimming pool is about 206 square feet. Estimate the radius of the pool, to the nearest foot. A. about 14 feet B. about 8 feet C. about 11 feet D. about 10 feet

OpenStudy (jdoe0001):

 \(\bf \textit{Area of a Circle} = \pi r^2\implies 206 = \pi r^2\) r = radius of the circle if you want to know what the radius is, solve for "r"